OpenStudy (anonymous):

find the product of 2 and 7

OpenStudy (saifoo.khan):

9

OpenStudy (anonymous):

14

OpenStudy (angela210793):

2*7=2+2+2+2+2+2+2=14

OpenStudy (saifoo.khan):

Oh! 14.

OpenStudy (saifoo.khan):

14

OpenStudy (saifoo.khan):

14

OpenStudy (anonymous):

product = multiplication

OpenStudy (anonymous):

thank you, English is not my first language, silly language to complicate things!

OpenStudy (saifoo.khan):

sorry, my bad.